# An Survey of the Evolution of Language Model-Based Dialogue Systems

### v1

@@ 6.2.3 Strategies for Mitigating Hallucinations
To mitigate hallucinations in language model-based dialogue systems, researchers have proposed several methods, which can be broadly categorized into five main approaches. One prominent approach is retrieval-augmentation [274], which involves incorporating information from external sources during the generation process to ensure the accuracy of the output. Model editing [275] is another approach adopted by researchers to mitigate hallucinations in language model-based dialogue systems. This involves precisely adjusting the model’s parameters to correct erroneous knowledge inherent in the system. Post-editing [276] is another commonly used approach to mitigate hallucinations in dialogue systems. After the model’s output is generated, post-editing techniques are applied to correct any detected hallucinations. This typically involves rule-based or pattern-based modifications to the generated output, ensuring that the final response is both coherent and accurate. Indeed, some methods aim to mitigate hallucinations by reducing the randomness in language model decoding. For instance, dynamic top-p sampling decoding [266] dynamically adjusts the randomness as the decoding process progresses, thus reducing the element of chance and helping to decrease the incidence of hallucinations. Lastly, better training strategies, such as RLHF [277, 263], have been developed to encourage the model to avoid making factually incorrect statements. By leveraging feedback from humans during the training process, these methods can guide the model towards generating more accurate and reliable responses.
@@ 6.3 Safety
Alongside the accelerated evolution of language models and dialogue systems, emerging safety concerns become increasingly crucial. Given the indispensable social impact of current LLM-based dialogue systems, their safety issues have garnered attention not only from the NLP research community [278, 279] but also from industry, philosophers, social scientists [280, 281], policymakers [282], etc. In this subsection, we begin by introducing the safety problem and its distinctive nature in dialogue systems which are directly interacting with human users. Subsequently, we offer a brief overview of existing safeguarding methods and discuss their limitations.
6.3.1 Understanding the Safety Issue
@@
The term safety encompasses a broad spectrum of issues [283], comprising, but not limited to, toxic language [284], social bias [285, 286], problematic moral or values inclination [287, 288], privacy leakage [289], and misleading or risky suggestions [290, 291]. It is noteworthy that the rigorous scope and taxonomy are still an ongoing discussion. Moreover, it is anticipatable that as the system’s capability and application scenarios broaden, the community may encounter novel safety challenges. For example, there is an emergent research trend of investigating prosocial dialogue systems [287], which are sensitive to social factors such as social norms [292, 293], moral stances [294, 288], and human values [295]. This progress stems from the dialogue systems’ expanding capabilities to engage in democratic conversations [280] that involve personal feelings, desires, and social relationships. In light of this, it is crucial to consistently evaluate the performance of dialogue systems to swiftly detect and tackle any new safety concerns that may arise.
6.3.2 Existing Safeguarding Methods
@@
Learning from enormous corpora from various sources, current dialogue systems inevitably learn undesired behavior. Mainstream safeguarding techniques can be categorized into two types: baked-in safety [296] and pipeline safety [278]. Baked-in safety refers to the methods that change the model at the training or decoding stage [297]. The majority of open-source LLMs have incorporated various safety training methods during the training or fine-tuning stages [298, 279], with the RLHF technique [4] being the most effective and widely-discussed option. This method can embed safety into models, however, it is not flexible nor explainable. Pipeline safety does not require modification to models, instead, it uses an add-on safety layer (usually a safety detector) to block unsafe input or output [299]. Nevertheless, this method cannot defend against the malicious use of open-sourced models. Another common and important obstacle in the above methods is the requirement of well-annotated corpora to train classifiers. Recently, there has been a novel research line utilizing the instruction following the ability of LLMs to perform self-critique [300] and self-correct [301]. The LLMs are instructed to serve as the classifier and make post-edition to its responses based on the classification results. Moreover, it is also instructed to self-redteaming [302] to explore high inductive user inputs and enhance the robustness automatically.
6.3.3 Challenges and Discussions
@@
The trade-off between safety and system performance has long been an intricate issue during the development of safe dialogue systems. [296] highlights the challenges of simultaneously maintaining safety and engagement in a dialogue system. Adding to this, Bai et al. [303] provides evidence that there exists a tension between ensuring harmlessness, i.e., safety, and maintaining helpfulness in such systems. This tension, further referred to by researchers as the alignment tax [304] remains an unresolved issue in the field. Another challenge is that the safety standards can change over time, culture, etc. [294] For example, mimic human was an important objective for old-days chatbots [3]. However, nowadays systems are trained to avoid anthropomorphism [305]. Furthermore, the criticism of generalizing the stances and values of a limited group of annotators has not been settled. Considering this, the methods that require a huge amount of training data and extensive training procedure is inflexible. A growing body of literature [288] advocates for steering clear of aligning AI models with specific values, since the human influence during the development lifecycle can inadvertently introduce systematic biases [306, 307]. Instead, scholars argue that shaping AI output through transparent and adaptable guidelines would be a more prudent approach [288]. However, the exploration of this direction remains in its infancy and warrants further investigation to develop a safe dialogue system.
7 Conclusion
This survey presents a comprehensive overview of the dialogue system’s evolution in tandem with advancements in language models, spanning from traditional Statistical Language Models (SLM) to Neural Language Models (NLM), Pre-trained Language Models (PLM), and culminating in the current state of Large Language Models (LLM). The exposition begins by establishing the contextual background of dialogue systems, delineating four distinct stages and a unified definition encompassing various dialogue types. Subsequently, we meticulously detail the progression and transformation of dialogue systems across different language model stages, emphasizing recent strides in LLM-based systems. Concurrently, we succinctly summarize key methodologies characterizing each modeling paradigm at every stage, providing a comprehensive snapshot of current advancements and guiding future directions. Our discourse extends to the application of large language models in addressing emerging challenges such as hallucinations and safety concerns, underscoring the relevance and adaptability of LLMs to evolving problem landscapes. This comprehensive survey not only encapsulates the historical trajectory but also offers insights into the current landscape and potential avenues for future research in dialogue systems.
Acknowledgments
This project is partially supported by HKSAR ITF Project ITT/008/22LP.

### v2

@@ 6.4.3. Strategies for Mitigating Hallucinations
To mitigate hallucinations in language model-based dialogue systems, researchers have proposed several methods, which can be broadly categorized into five main approaches. One prominent approach is retrieval-augmentation (Shuster et al., 2021), which involves incorporating information from external sources during the generation process to ensure the accuracy of the output. Model editing is another approach adopted by researchers to mitigate hallucinations in language model-based dialogue systems. This involves precisely adjusting the model’s parameters to correct erroneous knowledge inherent in the system. Post-editing (Dziri and et al., 2021) is another commonly used approach to mitigate hallucinations in dialogue systems. After the model’s output is generated, post-editing techniques are applied to correct any detected hallucinations. This typically involves rule-based or pattern-based modifications to the generated output, ensuring that the final response is both coherent and accurate. Indeed, some methods aim to mitigate hallucinations by reducing the randomness in language model decoding. For instance, dynamic top-p sampling decoding (Lee et al., 2023) dynamically adjusts the randomness as the decoding process progresses, thus reducing the element of chance and helping to decrease the incidence of hallucinations. Lastly, better training strategies, such as RLHF (Xue and et al., 2023), have been developed to encourage the model to avoid making factually incorrect statements. By leveraging feedback from humans during the training process, these methods can guide the model towards generating more accurate and reliable responses.
@@ 6.5. Safety
Alongside the accelerated evolution of language models and dialogue systems, emerging safety concerns become increasingly crucial. Given the indispensable social impact of current LLM-based dialogue systems, their safety issues have garnered attention not only from the NLP research community (Deng and et al., 2023a; Bai et al., 2022) but also from industry, philosophers, social scientists (Irving and Askell, 2019), policymakers (Zeng et al., 2018), etc. In this subsection, we begin by introducing the safety problem and its distinctive nature in dialogue systems which are directly interacting with human users. Subsequently, we offer a brief overview of existing safeguarding methods and discuss their limitations.
6.5.1. Understanding the Safety Issue
@@
The term safety encompasses a broad spectrum of issues (Weidinger et al., 2022), comprising, but not limited to, toxic language (et al., 2022a), social bias (et al., 2020a; Zhou and et al., 2022), problematic moral or values inclination (et al., 2022c), and misleading or risky suggestions (Mei et al., 2022; Sun et al., 2022). It is noteworthy that the rigorous scope and taxonomy are still an ongoing discussion. Moreover, it is anticipatable that as the system’s capability and application scenarios broaden, the community may encounter novel safety challenges. For example, there is an emergent research trend of investigating prosocial dialogue systems (et al., 2022c), which are sensitive to social factors such as social norms (et al., 2022d), moral stances (et al., 2023c), and human values (et al., 2023b). This progress stems from the dialogue systems’ expanding capabilities to engage in democratic conversations that involve personal feelings, desires, and social relationships. In light of this, it is crucial to consistently evaluate the performance of dialogue systems to swiftly detect and tackle any new safety concerns that may arise.
6.5.2. Existing Safeguarding Methods
@@
Learning from enormous corpora from various sources, current dialogue systems inevitably learn undesired behavior. Mainstream safeguarding techniques can be categorized into two types: baked-in safety (et al., 2020b) and pipeline safety (Deng and et al., 2023a). Baked-in safety refers to the methods that change the model at the training or decoding stage (et al., 2021b). The majority of open-source LLMs have incorporated various safety training methods during the training or fine-tuning stages (Thoppilan et al., 2022; Bai et al., 2022), with the RLHF technique (Ouyang et al., 2022) being the most effective and widely-discussed option. This method can embed safety into models, however, it is not flexible nor explainable. Pipeline safety does not require modification to models, instead, it uses an add-on safety layer (usually a safety detector) to block unsafe input or output (Dinan and et al., 2022). Nevertheless, this method cannot defend against the malicious use of open-sourced models. Another common and important obstacle in the above methods is the requirement of well-annotated corpora to train classifiers. Recently, there has been a novel research line utilizing the instruction following the ability of LLMs to perform self-critique (Wang et al., 2023c) and self-correct (et al., 2023a). The LLMs are instructed to serve as the classifier and make post-edition to its responses based on the classification results. Moreover, it is also instructed to self-redteaming (Wang et al., 2023d) to explore high inductive user inputs and enhance the robustness automatically.
6.5.3. Challenges and Discussions
@@
The trade-off between safety and system performance has long been an intricate issue during the development of safe dialogue systems. (et al., 2020b) highlights the challenges of simultaneously maintaining safety and engagement in a dialogue system. Adding to this, Bai et al. (Bai and et al., 2022) provides evidence that there exists a tension between ensuring harmlessness, i.e., safety, and maintaining helpfulness in such systems. This tension, further referred to by researchers as the alignment tax (et al., 2021a) remains an unresolved issue in the field. Another challenge is that the safety standards can change over time, culture, etc. (et al., 2023c) For example, mimic human was an important objective for old-days chatbots (Zhou et al., 2020). However, nowadays systems are trained to avoid anthropomorphism (Glaese and et al., 2022). Furthermore, the criticism of generalizing the stances and values of a limited group of annotators has not been settled. Considering this, the methods that require a huge amount of training data and extensive training procedure is inflexible. A growing body of literature (Zhou and et al., 2024) advocates for steering clear of aligning AI models with specific values, since the human influence during the development lifecycle can inadvertently introduce systematic biases (Sap and et al., 2022; et al., 2022b). Instead, scholars argue that shaping AI output through transparent and adaptable guidelines would be a more prudent approach (Zhou and et al., 2024). However, the exploration of this direction remains in its infancy and warrants further investigation to develop a safe dialogue system.
7. Conclusion
This survey provides a comprehensive overview of the evolution of dialogue systems alongside advances in language modeling-from Statistical Language Models (SLMs) to Neural Language Models (NLMs), Pre-trained Language Models (PLMs), and the current generation of Large Language Models (LLMs). We begin by outlining the contextual background of dialogue systems, defining four developmental stages to trace the progression of dialogue systems, highlighting clear bottom-up evolution and clear path for future direction. Our discourse extends to the challenges and applications of LLM-based dialogue system in real-world tasks.